Hello, Please help me solve these
(a) The probability of a bank customer defaulting on a loan is 0.04 if the economy is doing well (high) and 0.13 if the economy is not high. Suppose the probability that the economy is high is 0.65, what is the probability the customer will default on the loan?
(b) The probability that a graduating student will get an offer for job A is 0.35 and the probability that this student will get an offer for job B is 0.20. If the probability of this student getting both job offers is 0.05, what is the probability that this student will get at least one of the two job offers?

Answer by ikleyn(53765)

;
(a) The probability of a bank customer defaulting on a loan is 0.04 if the economy is doing well (high) and 0.13
if the economy is not high. Suppose the probability that the economy is high is 0.65, what is the probability
the customer will default on the loan?
(b) The probability that a graduating student will get an offer for job A is 0.35 and the probability that this student
will get an offer for job B is 0.20. If the probability of this student getting both job offers is 0.05, what is the probability
that this student will get at least one of the two job offers?
~~~~~~~~~~~~~~~~~

(a)  P = 0.65*0.04 + (1-0.65)*0.13.



(b)  P(at least one) = P(A or B) = P(A) + P(B) - P(A and B) = 0.35 + 0.2 - 0.05.